10.20) The law practice of Smyth, Pritcher and Dunne has monthly fixed costs of $115000, EBIT of $289000 and depreciation charges on its office furniture and computers of $5800. Calculate the cash flow DOL for this company.

10.29) Twilight Candles will be producing a new line of dripless candles. If the large factory is chosen, the cost per unit to produce each candle will be $2.48, while the cost per unit will be $6.84 for the small factory. The large factory would have fixed costs of $2.4 milion and depreciation expense of $300000 per year, while those expense would be $52000 and $100000 in the small factory. Calculate the accounting operating profit break-even point for both factory choices for Twilight Candles.

10.30) Swiffer Cleaning Solutions will be producing a new line of brooms. If the large factory is chosen, the cost per unit to produce each broom will be $2.84, while the cost per unit will be $6.67 for the small factory. The large factory would have fized cash costs of $2.2million and a depreciation expense of $300000 per year, while those expenses would be $460000 and $100000 in the small factory. Calculate the number of brooms for Swiffer Cleaning Solutions for which the accounting operating profit is the same, regardless of the factory choice.

10.31)Aromatic Ltd will be producing a new line of essential oils. If the large factory chosen, the cost per unit to produce each oil will be $2.50, while the cost per unit will be $6.78 for the small factory. The large factory would have fixed cash costs of $2 million and a depreciation expense of $300000 per year, while those expenses would be $500000 and $100000 in the small factory. Calculate the pre-tax operating cash flow break-even point for both factory choices for Aromatic Ltd.

10.40)Tim’s Orange Juice is starting to develop a new product for which the fixed cash expenditures are expected to be $93000. The projected EBIT is $132000 and Accounting DOL will be 2.2. What is the Cash flow DOL for the company?
Answer #1

10.20

Cash flow DOL = 1 + Fixed costs / EBITDA = 1 + 115,000 / 289,000 + 5800 = 1.39

10.29

Small factory Large Factory
Selling price [ assumed each is sold for $10] 10 10.00
Less: Cost per unit 6.84 2.48
Contribution per unit 3.16 7.52
Fixed costs 152000 2700000
Break even units
[Fixed costs/ Contribution per unit]
48101 359043

10.30 . Cross over level of sales = [(FC + Depreciation ) Alt 1 + (FC + Depreciation ) Alt 1] / (Contribution per unit Alt 1 - Contribution per unit Alt 2)

= ( 2,500,000 + 560,000) / [ (10- 2.84) - (10- 6.67)]

= 506,527 units

10. 31 . Cash flow break even point = Fixed costs/ Contribution per unit

Large factory = 2,000,00/ (10 2.5) = 266,667 units

Small factory= 500,000/ (10 6.78) = 155,280 units

10.40. Accounting DOL = 1 + (Fixed costs +depreciation)/ EBIT

2.2 = 1+ (93000 +Depreciation) / 132000

Depreciation = 65400

Cash flow DOL = 1 + Fixed costs / EBITDA = 1+ 93000 / ( 132000 + 65400) = 1.47